OpenStudy (kohai):

A rock is dropped from rest from a height h above the ground. It falls and hits the ground with a speed of 11m/s. From what height should the rock be dropped so that its speed on hitting the ground is 22m/s? Neglect air resistance.

OpenStudy (zephyr141):

not too sure if i did it right, but i used V^2=V(0)^2+2aDelta(y). V=22m/s V(0)=0m/s a=g=-9.8 or -9.81 and Delta(y)=h and solved for h. not sure if it's correct but try it out. hopefully someone more seasoned answers your question.

OpenStudy (kohai):

I think it's 4h

OpenStudy (zephyr141):

oh. where did the 4 come from? i'm having trouble in the same class as well.

OpenStudy (kohai):

 I got the beginning from a few formulas